Explain how reducing the budget deficit would impact the trade deficit. Using macroeconomic equations, how could reducing the budget deficit decrease the trade deficit subsequently?
Answer #1

Budget deficit and trade deficit are closely related , change in budget deficit can have impact on trade deficit. If we consider the national income accounting model of macroeconomics where Y=C+I+G+X-M where Y= national income or GDP, C= consumption, I investment, G= government spending, X-M= net export.

another macroeconomic equation of calculating GDP is; Y= C+S+T where S= savings and T = tax

taking both the equation in consideration, S= G+NX++I-T where NX= X-M

now, (S-I)+(T-G)=NX

if (T-G) is negative we have budget deficit.assuming the economy is at potential output already which means GDP is fixed, in this case if budget deficit reduces keeping saving constant either investment must rise or net export must rise thus trade deficit will reduce susequently.